# Optimising Communication Control Factors for Energy Consumption in Rural V2X

### v2

@@ Abstract
Connected braking can reduce fatal collisions in connected and autonomous vehicles (CAVs) by using reliable, low-latency 5G New Radio (NR) links, especially NR Sidelink Vehicle-to-Everything (V2X). In rural areas, road side units are sparse and power-constrained or off-grid, so energy efficiency must be considered alongside safety. This paper studies how three communication control factors including subcarrier spacing ($\mathrm{SCS}$), modulation and coding scheme ($\mathrm{MCS}$), and transmit power ($P_{\mathrm{t}}$) should be configured to balance safety and energy consumption in rural scenarios in light and heavy traffic scenarios. Safety is quantified by the packet receive ratio ($\mathrm{PRR}$) against the minimum communication distance $D_{\mathrm{comm}}$, defined as the distance that the vehicle travels during the transmission of the safety message. Results show that, under heavy traffic, increasing $P_{\mathrm{t}}$ and selecting a low-rate $\mathrm{MCS}$ at $\mathrm{SCS}=30$ kHz sustains high $\mathrm{PRR}$ at $D_{\mathrm{comm}}$, albeit with higher energy cost. In light traffic, maintaining lower $P_{\mathrm{t}}$ with low $\mathrm{MCS}$ levels achieves a favorable reliability-energy trade-off while preserving acceptable $\mathrm{PRR}$ at $D_{\mathrm{comm}}$. These findings demonstrate the necessity of adaptive, energy-aware strategy to guarantee both safety and energy efficiency in rural V2X systems.
@@ I Introduction
Connected and autonomous vehicle (CAV) technologies are advancing rapidly in urban areas, whilst rural deployment continues to lag behind. One of the primary barriers is the limited interest from network operators and policy makers in investing in dedicated infrastructure for sparsely populated regions, where naturally, the return on investment is often less attractive. However, rural road users expect the same level of connectivity safety as those in urban settings [Scenario]. This growing expectation raises an important research question about how to ensure a reliable and robust level of CAV safety in rural areas, where communication resources are limited. Rather than relying on the expansion of infrastructure from urban to rural, it becomes critical to explore how to maximize the efficiency of available rural resources to support Vehicle-to-Everything (V2X)-based safety systems in rural driving scenarios. In this context, it is vital to address the gap by exploring strategies for a resource-efficient V2X safety deployment, which is specifically tailored to the unique challenges of rural environments.
@@
Typically, CAVs rely on integrated sensors such as cameras, radar, and lidar to process braking decisions by inferring data from their surroundings. However, in complex scenarios such as roundabouts, road merges, and junctions, sensor accuracy can degrade due to obstructions or observation failures, exposing a key limitation of sensor-based systems [Sensorfusion]. Studies show that safe braking distances in such environments are greatly improved through connectivity, which allows vehicles and drivers to be alerted to potential hidden hazards in advance [Jason]. Cooperative V2X communication, including Vehicle-to-Vehicle (V2V) and Vehicle-to-Infrastructure (V2I), extends environmental awareness and enables earlier, safer braking decisions [Humanjiangbrake]. Among the available V2X technologies, DSRC-based V2V and 5G NR V2X are the most widely adopted. Compared to DSRC V2V, 5G New-Radio (NR) V2X provides enhanced performance, offering higher data rates, lower communication latency, greater scalability, and more robust support for high-mobility traffic scenarios [5GVSDSRCHOLOGRAMS]. These advantages enable 5G NR V2X to better meet the stringent safety requirements of CAV systems [ToDSRCor5G]. Across all 5G NR V2X technologies, Sidelink V2X offers a more reliable alternative by supporting direct, low-latency vehicle communication without relying on fixed infrastructure [sidelinkv2v]. This makes Sidelink V2X a robust solution for maintaining connectivity and enhancing reaction times, even in sparse or infrastructure-limited environments [sidelinkv2x].
@@
While V2X communication is primarily designed to improve traffic efficiency and road safety, the energy cost associated with maintaining reliable wireless links becomes especially critical in rural deployments. In such environments, roadside infrastructure is often sparse and may rely on limited or off-grid power sources. Unlike sensing or control operations that are event-driven, wireless communication requires continuous use of energy for data transmission and reception. Therefore, reducing the communication energy is essential to ensure long-term system sustainability and cost-effectiveness in rural areas. Communication control factors, including subcarrier spacing ($\mathrm{SCS}$), transmit power, and modulation and coding scheme ($\mathrm{MCS}$) play a direct role in balancing energy consumption and communication reliability. Wider $\mathrm{SCS}$ and higher $\mathrm{MCS}$ values enable higher data throughput and can reduce transmission time, which lowers energy consumption per bit when the channel is stable [SCSMCS]. However, these configurations are less robust in the presence of interference or fading, leading to potential retransmissions and energy waste, especially when link quality is poor. Increasing transmit power improves signal robustness, but adds to the overall energy cost, which is particularly undesirable in power-constrained rural infrastructure. When it comes to traffic conditions, they are typically classified as “light” or “heavy”. Light traffic, common on rural roads, often features low vehicle density and higher average speeds, where wider $\mathrm{SCS}$ and higher $\mathrm{MCS}$ can be exploited for efficient communication. In contrast, temporary congestion or mixed traffic conditions may demand more conservative configurations to preserve reliability, especially under variable channel conditions. Given the unique challenges of rural V2X communication, such as limited infrastructure, dynamic vehicle movement, and limited energy resources, it is essential to systematically analyse how communication control settings affect both energy efficiency and safety-critical performance. Such analysis supports the design of adaptive, energy-aware communication strategies that ensure reliable operation without overprovisioning power or compromising safety.
@@
In this paper, building on our previous work [ZhanleIVs], we extend the investigation to more complex rural traffic scenarios. First, we analyse minimum safety-critical distances that the ego vehicle travels while the braking signal is being transmitted. This analysis is conducted under varying rural traffic conditions to characterise the system’s behaviour in dynamic environments. Second, by evaluating the packet reception rate ($\mathrm{PRR}$) against minimum safety-critical distances, we investigate the optimisation of key communication control factors which are $\mathrm{SCS}$, $\mathrm{MCS}$, and the transmit power of the road side unit ($P_{\mathrm{t}}$). This analysis aims to identify effective parameter combinations that can ensure reliable communication and maintain a safe critical distance under diverse traffic conditions. Finally, we conduct an energy consumption analysis under both light and heavy traffic scenarios to determine energy-efficient communication strategies that maintain safety-critical performance while minimising energy cost.
@@ I-A Related Research
Connected braking is considered as a safety feature used for assisting a driver, a remote operator, or an autonomous controller to stop the vehicle in advance through V2X communications. In our previous work [ZhanleIVs], we developed a metric, so-called $D_{\mathrm{comm}}$, as the minimum safe critical distance that the ego vehicle traveled during the time that braking signals are transmitted. It considers a scenario that the ego vehicle, due to its sensors, is unaware of the sudden braking action of a vehicle ahead in a rural setting. As a result, a crash is imminent unless an avoidance maneuver or braking action is executed. In the context of connected braking, the ego vehicle communicates with the target vehicle or the RSU and triggers the brake based on delivered data packets. A larger packet delay leads to a slower connected braking, which increases $D_{\mathrm{comm}}$ accordingly. In general, the increase of $D_{\mathrm{comm}}$ is mainly caused by the increase of packet error rate ($\mathrm{PER}$), which is defined as the total number of received error packets against the total number of packet transmitted, and it is the main indicator of V2X propagation delay [PERinWCL]. In [ZhanleIVs], we investigate the impact of transmit power on the V2V $\mathrm{PER}$ in a DSRC channel. An analysis on $D_{\mathrm{comm}}$ is provided to demonstrate connected braking requirements within a $100$ $\mathrm{m}$ in a rural road-following scenario.
@@
The optimisation of communication control factors in V2X communication remains an open research challenge [v2xchallenge]. In practice, the increasing complexity of road environments, driven by a higher number of road users, additional lanes, extended road lengths, and higher vehicle speeds make it even more difficult to maintain the required levels of safety and reliability. This challenge is closely linked to the allocation of communication control factors, such as transmit power, $\mathrm{SCS}$, and $\mathrm{MCS}$, which must be dynamically adjusted to match traffic conditions while avoiding excessive energy consumption. As the number of vehicles participating in the V2X network increases, the communication system must handle significantly higher data volumes, which can lead to network congestion, packet loss, and increased message delivery delays [DSRCplatooning]. These conditions directly impact the reliability of V2X connections, especially in environments with a lack of communication resource such as rural. In parallel, higher initial vehicle speeds further complicate communication demands. The authors of [V2Xbraking] show that increased speeds require faster reaction times, which in turn demand a higher signal-to-noise ratio (SNR) to maintain low latency performance and increase the required $D_{\mathrm{comm}}$. This adds pressure on the critical-safety system to deliver Basic Safety Messages (BSMs) with minimal delay [BSMs]. Collectively, these factors increase the likelihood of missed or delayed hazard alerts in complex, high-mobility scenarios, making the real-time monitoring and control of $D_{\mathrm{comm}}$ more challenging. These challenges highlight the need for a proper setting of communication factors that can ensure timely and reliable message delivery under varying traffic and mobility conditions.
@@
The impact of communication control factors on V2X communication has been investigated in a number of studies [V2XRA-Pt, V2XRA-SCS, V2XRA3, V2Xsurvey, V2XRA-MCS-NR1, V2XRA-MCS-NR2]. The authors of [V2XRA-Pt] have investigated the impact of ${P}_{\mathrm{t}}$ on communication quality in urban scenario. It shows that a higher ${P}_{\mathrm{t}}$ improves the received SNR, especially over long distances or in high interference environments. The research proposed in [V2XRA-SCS] has specifically analysed the optimisation strategy of $\mathrm{SCS}$ in V2X communication systems. Using a relatively high $\mathrm{SCS}$, such as $30$ $\mathrm{kHz}$, shortens the duration of each symbol, making it particularly suitable for low-latency applications. The study introduced in [V2XRA3, V2Xsurvey] analyses the optimisation of both $\mathrm{SCS}$ and $\mathrm{MCS}$ in urban V2X scenarios. Their findings indicate that a proper tuning of $\mathrm{SCS}$ and $\mathrm{MCS}$ enables the system to better adapt to varying traffic loads and channel conditions, thereby supporting more responsive and reliable data exchange, particularly under high vehicular densities. The research [V2XRA-MCS-NR1, V2XRA-MCS-NR2] focus on $\mathrm{MCS}$ optimisation within urban 5G NR V2X scenarios. It shows that appropriately adapting the $\mathrm{MCS}$ to match prevailing channel conditions can significantly reduce PER. Despite providing useful insights, the above works [V2XRA-Pt, V2XRA-SCS, V2XRA1, V2XRA-MCS, V2XRA-MCS-NR1, V2XRA-MCS-NR2] do not address a joint optimisation of ${P}_{\mathrm{t}}$, $\mathrm{SCS}$, and $\mathrm{MCS}$. This gap is especially notable in the context of 5G NR V2X for rural traffic scenarios, where the unique channel characteristics demand a holistic approach to communication control. Another notable gap in [V2XRA-Pt, V2XRA-SCS, V2XRA1, V2XRA-MCS, V2XRA-MCS-NR1, V2XRA-MCS-NR2] is the lack of consideration of energy consumption in the optimisation of communication control factors. Although previous works have focused on improving reliability and latency, it has largely overlooked the trade-offs between performance and energy efficiency, which is an increasingly important concern, particularly in resource-constrained V2X deployments such as rural traffic scenarios.
@@
Several studies have considered the energy aspect of V2X communication, recognising its importance in sustaining large-scale, real-time vehicle connectivity [energy1, V2XRA-Pt2, EnergySCS, EnergySCS2, EnergyMCS]. These works typically explore how different communication control factors can be optimised to reduce energy consumption without compromising reliability. The energy-efficient resource allocation problem in urban V2X scenarios is considered in [energy1, V2XRA-Pt2]. It indicates that the appropriate control of ${P}_{\mathrm{t}}$ and interference management are important for energy-efficient V2X. In addition, $\mathrm{SCS}$ and $\mathrm{MCS}$ can also come at the cost of increased power consumption. The research conducted in [EnergySCS] shows that higher $\mathrm{SCS}$, such as $30$ $\mathrm{kHz}$, can lead to higher energy requirements for accurate transmission and reception due to increased signal bandwidth and processing demands, especially under different traffic conditions. In contrast, it was shown that such low-priority and delay-tolerant applications as traffic monitoring or non-time-sensitive updates, benefit from lower $\mathrm{SCS}$ such as $15$ $\mathrm{kHz}$ [EnergySCS2]. Moreover, lower $\mathrm{MCS}$ levels can conserve energy and minimise resource usage. As demonstrated in [EnergyMCS], selecting a lower $\mathrm{MCS}$ in challenging environments leads to improved transmission efficiency and reduced energy consumption without significantly compromising throughput. However, these studies [energy1, V2XRA-Pt2, EnergySCS, EnergySCS2, EnergyMCS] often overlook rural environments, where vehicles are typically more spaced out and signals travel over longer distances with fewer obstructions. Such conditions demand a joint optimisation of all three communication control factors to ensure reliable communication while maintaining energy efficiency.
@@
Motivated by these research gaps, this paper investigates the impact of three communication control factors ${P}_{\mathrm{t}}$, $\mathrm{SCS}$, and $\mathrm{MCS}$ on energy consumption and communication performance in different rural V2X scenarios. Performance is evaluated using $\mathrm{PRR}$ and the $D_{\mathrm{comm}}$, reflecting the ability to deliver safety messages within safety-critical distances. By analysing $\mathrm{PRR}$ and energy consumption under different traffic densities and vehicle speeds, the study aims to identify optimal configurations that ensure reliable, low-latency communication while minimising energy expenditure.
I-B Main Contributions
The main contributions of this paper now can be summarised as follows. Firstly, we investigate the impact of three key communication control factors including $\mathrm{SCS}$, transmit power of RSU, and $\mathrm{MCS}$ on safety and energy consumption in a connected braking scenario. Our rural scenario represents random layouts in light and heavy traffic conditions, where CAVs exchange time-critical braking messages via 5G NR Sidelink V2X communication. Our second contribution is the development of analytical formulations for the $D_{\mathrm{comm}}$, $\mathrm{PRR}$ and energy consumption, which together describe the achievable braking reliability over distance. These formulations provide a comprehensive analysis, linking communication control factors to safety-critical performance indicators. Finally, the impact of communication control factors on the safety and energy consumption is validated through extensive simulations, offering valuable insights into the trade-offs between safety and energy consumption. Using the analytical results, we identify the appropriate configurations of $\mathrm{SCS}$, ${P}_{\mathrm{t}}$, and $\mathrm{MCS}$ that jointly optimise safety and energy performance. The analysis forms a baseline for adaptive control strategies on them in future rural CAV safety applications.
The remainder of this paper is organised as follows. Section II introduces the system model and the simulation setup. In Section III, we formulate $D_{\mathrm{comm}}$ and energy consumption. In Section IV, we illustrate the simulation results and explore the impact on $\mathrm{PRR}$, $D_{\mathrm{comm}}$ from tuning communication control factors. Section V provides quantitative analysis on the results, and impact of communication control factors on safety and energy consumption. Finally, Section VI summarises the main findings of this paper.

### v3

# Toward Safe and Energy-Efficient 5G NR V2X Communications in Rural Environments

@@ Abstract
Connected braking can reduce fatal collisions in connected and autonomous vehicles (CAVs) by using reliable, low-latency 5G New Radio (NR) links, especially NR Sidelink Vehicle-to-Everything (V2X). In rural areas, road side units are sparse and power-constrained, so energy efficiency must be considered alongside safety. This paper studies how three communication control factors including subcarrier spacing ($\mathrm{SCS}$), modulation and coding scheme ($\mathrm{MCS}$), and transmit power ($P_{\mathrm{t}}$) should be configured to balance safety and energy consumption in rural scenarios in light and heavy traffic scenarios. Safety is quantified by the packet receive ratio ($\mathrm{PRR}$) against the minimum communication distance $D_{\mathrm{comm}}$, defined as the distance that the vehicle travels during the transmission of the safety message. Results show that, under heavy traffic, increasing $P_{\mathrm{t}}$ and selecting a low-rate $\mathrm{MCS}$ at $\mathrm{SCS}=30$ kHz sustains high $\mathrm{PRR}$ at $D_{\mathrm{comm}}$, albeit with higher energy cost. In light traffic, maintaining lower $P_{\mathrm{t}}$ with low $\mathrm{MCS}$ levels achieves a favorable reliability-energy trade-off while preserving acceptable $\mathrm{PRR}$ at $D_{\mathrm{comm}}$. These findings demonstrate the necessity of adaptive, energy-aware strategy to guarantee both safety and energy efficiency in rural V2X systems.
@@ I Introduction
Connected and autonomous vehicle (CAV) technologies are advancing rapidly in urban areas, whilst rural deployment continues to lag behind. One of the primary barriers is the limited interest from network operators and policy makers in investing in dedicated infrastructure for sparsely populated regions, where naturally, the return on investment is often less attractive. However, rural road users expect the same level of connectivity safety as those in urban settings [25]. This growing expectation raises an important research question about how to ensure a reliable and robust level of CAV safety in rural areas, where communication resources are limited. Rather than relying on the expansion of infrastructure from urban to rural, it becomes critical to explore how to maximize the efficiency of available rural resources to support Vehicle-to-Everything (V2X)-based safety systems in rural driving scenarios. In this context, it is vital to address the gap by exploring strategies for a resource-efficient V2X safety deployment, which is specifically tailored to the unique challenges of rural environments.
@@
Typically, CAVs rely on integrated sensors such as cameras, radar, and lidar to process braking decisions by inferring data from their surroundings. However, in complex scenarios such as roundabouts, road merges, and junctions, sensor accuracy can degrade due to obstructions or observation failures, exposing a key limitation of sensor-based systems [30]. Studies show that safe braking distances in such environments are greatly improved through connectivity, which allows vehicles and drivers to be alerted to potential hidden hazards in advance [32]. Cooperative V2X communication, including Vehicle-to-Vehicle (V2V) and Vehicle-to-Infrastructure (V2I), extends environmental awareness and enables earlier, safer braking decisions [10]. Among the available V2X technologies, DSRC-based V2V and 5G NR V2X are the most widely adopted. Compared to DSRC V2V, 5G New-Radio (NR) V2X provides enhanced performance, offering higher data rates, lower communication latency, greater scalability, and more robust support for high-mobility traffic scenarios [22]. These advantages enable 5G NR V2X to better meet the stringent safety requirements of CAV systems [7]. Across all 5G NR V2X technologies, Sidelink V2X offers a more reliable alternative by supporting direct, low-latency vehicle communication without relying on fixed infrastructure [14]. This makes Sidelink V2X a robust solution for maintaining connectivity and enhancing reaction times, even in sparse or infrastructure-limited environments [21].
@@
While V2X communication is primarily designed to improve traffic efficiency and road safety, the energy cost associated with V2X infrastructure has been considered from multiple perspectives, including sensing and perception workloads, onboard computing, and system-level power management strategies [16, 34]. Among these components, the radio frequency (RF) communication subsystem remains an important and controllable contributor to energy consumption, as transmission power, and retransmissions are directly influenced by communication configurations, channel and traffic conditions [29, 27]. The energy cost associated with maintaining reliable wireless links becomes especially critical in rural deployments. In such environments, road side infrastructure is often sparse and may rely on limited power sources. Unlike sensing or control operations that are event-driven, wireless communication requires continuous use of energy for data transmission and reception. Therefore, reducing the RF energy is essential to ensure long-term system sustainability and cost-effectiveness in rural areas. Communication control factors, including subcarrier spacing ($\mathrm{SCS}$), transmit power, and modulation and coding scheme ($\mathrm{MCS}$) play a direct role in balancing RF energy consumption and communication reliability. Wider $\mathrm{SCS}$ and higher $\mathrm{MCS}$ values enable higher data throughput and can reduce transmission time, which lowers RF energy consumption per bit when the channel is stable [6]. However, these configurations are less robust in the presence of interference or fading, leading to potential retransmissions and energy waste, especially when link quality is poor. Increasing transmit power improves signal robustness, but adds to the overall energy cost, which is particularly undesirable in power-constrained rural infrastructure. When it comes to traffic conditions, they are typically classified as “light” or “heavy”. Light traffic, common on rural roads, often features low vehicle density and higher average speeds, where wider $\mathrm{SCS}$ and higher $\mathrm{MCS}$ can be exploited for efficient communication. In contrast, temporary congestion or mixed traffic conditions may demand more conservative configurations to preserve reliability, especially under variable channel conditions. Given the unique challenges of rural V2X communication, such as limited infrastructure, dynamic vehicle movement, and limited energy resources, it is essential to systematically analyse how communication control settings affect both energy efficiency and safety-critical performance. Such analysis supports the design of adaptive, energy-aware communication strategies that ensure reliable operation without overprovisioning power or compromising safety.
@@
In this paper, building on our previous work [33], we extend the investigation to more complex rural traffic scenarios. First, we analyse minimum safety-critical distances that the ego vehicle travels while the braking signal is being transmitted. This analysis is conducted under varying rural traffic conditions to characterise the system’s behaviour in dynamic environments. Second, by evaluating the packet reception rate ($\mathrm{PRR}$) with respect to minimum safety-critical distances, we analyse the impact of $\mathrm{SCS}$, $\mathrm{MCS}$, and RSU transmit power ($P_{\mathrm{t}}$) on safety-relevant reliability. This analysis performs design-space exploration over predefined 5G NR communication control configurations to identify parameter combinations that support reliable communication and maintain a safe critical distance under diverse rural V2X traffic conditions. Finally, we conduct an energy consumption analysis under both light and heavy traffic scenarios to determine energy-efficient communication strategies that maintain safety-critical performance while minimising energy cost, and provide guidance on configuration selection based on observed safety performance to energy trade-offs.
@@ I-A Related Research
Connected braking is considered as a safety feature used for assisting a driver, a remote operator, or an autonomous controller to stop the vehicle in advance through V2X communications. From the communication perspective, latency is a key safety metric, as it determines how quickly safety-relevant information is delivered. In practice, packet reliability and retransmissions further affect whether delivery occurs within the required time window. From the vehicle perspective, speed is particularly critical in rural environments, as it directly determines braking distance and the required safety margin. In our previous work [33], we developed a metric, so-called $D_{\mathrm{comm}}$, as the minimum safe critical distance that the ego vehicle traveled during the time that braking signals are transmitted. It considers a scenario that the ego vehicle, due to its sensors, is unaware of the sudden braking action of a vehicle ahead in a rural setting. As a result, a crash is imminent unless an avoidance maneuver or braking action is executed. The metric $D_{\mathrm{comm}}$ implicitly incorporates these factors by representing the maximum distance over which safety-relevant information can be exchanged with sufficient reliability, thereby mapping communication performance into a spatial safety margin that is directly interpretable in terms of vehicle dynamics and safety requirements. In the context of connected braking, the ego vehicle communicates with the target vehicle or the RSU and triggers the brake based on delivered data packets. A larger packet delay leads to a slower connected braking, which increases $D_{\mathrm{comm}}$ accordingly. In general, the increase of $D_{\mathrm{comm}}$ is mainly caused by the increase of packet error rate ($\mathrm{PER}$), which is defined as the total number of received error packets against the total number of packet transmitted, and it is the main indicator of V2X propagation delay [31]. In [33], we investigate the impact of transmit power on the V2V $\mathrm{PER}$ in a DSRC channel. An analysis on $D_{\mathrm{comm}}$ is provided to demonstrate connected braking requirements within a $100$ $\mathrm{m}$ in a rural road-following scenario.
@@
The performance evaluation of communication control factors in V2X communication remains an open research challenge [17]. In practice, the increasing complexity of road environments, driven by a higher number of road users, additional lanes, extended road lengths, and higher vehicle speeds make it even more difficult to maintain the required levels of safety and reliability. This challenge is closely linked to the allocation of communication control factors, such as transmit power, $\mathrm{SCS}$, and $\mathrm{MCS}$, which must be dynamically adjusted to match traffic conditions while avoiding excessive energy consumption. As the number of vehicles participating in the V2X network increases, the communication system must handle significantly higher data volumes, which can lead to network congestion, packet loss, and increased message delivery delays [13]. These conditions directly impact the reliability of V2X connections, especially in environments with a lack of communication resource such as rural. In parallel, higher initial vehicle speeds further complicate communication demands. The authors of [36] show that increased speeds require faster reaction times, which in turn demand a higher signal-to-noise ratio (SNR) to maintain low latency performance and increase the required $D_{\mathrm{comm}}$. This adds pressure on the critical-safety system to deliver Basic Safety Messages (BSMs) with minimal delay [20]. Collectively, these factors increase the likelihood of missed or delayed hazard alerts in complex, high-mobility scenarios, making the real-time monitoring and control of $D_{\mathrm{comm}}$ more challenging. These challenges highlight the need for a proper setting of communication factors that can ensure timely and reliable message delivery under varying traffic and mobility conditions.
@@
The impact of communication control factors on V2X communication has been investigated in a number of studies [12, 24, 18, 9, 15, 28]. The authors of [12] have investigated the impact of ${P}_{\mathrm{t}}$ on communication quality in urban scenario. It shows that a higher ${P}_{\mathrm{t}}$ improves the received SNR, especially over long distances or in high interference environments. The research proposed in [24] specifically evaluated the impact of $\mathrm{SCS}$ on V2X communication performance. Using a relatively high $\mathrm{SCS}$, such as $30$ $\mathrm{kHz}$, shortens the duration of each symbol, making it particularly suitable for low-latency applications. The study introduced in [18, 9] evaluates V2X communication performance under different $\mathrm{SCS}$ and $\mathrm{MCS}$ configurations in urban scenarios. Their findings indicate that a proper tuning of $\mathrm{SCS}$ and $\mathrm{MCS}$ enables the system to better adapt to varying traffic loads and channel conditions, thereby supporting more responsive and reliable data exchange, particularly under high vehicular densities. The research [15, 28] focus on $\mathrm{MCS}$ within urban 5G NR V2X scenarios. It shows that appropriately adapting the $\mathrm{MCS}$ to match prevailing channel conditions can significantly reduce PER. Despite providing useful insights, the above works [12, 24, 18, 9, 15, 28] do not address a joint performance evaluation of ${P}_{\mathrm{t}}$, $\mathrm{SCS}$, and $\mathrm{MCS}$. This gap is especially notable in the context of 5G NR V2X for rural traffic scenarios, where the unique channel characteristics demand a holistic approach to communication control. Another notable gap in [12, 24, 18, 9, 15, 28] is the lack of consideration of RF energy consumption in the performance evaluation of communication control factors. Although previous works have focused on improving reliability and latency, it has largely overlooked the trade-offs between performance and energy efficiency, which is an increasingly important concern, particularly in resource-constrained V2X deployments such as rural traffic scenarios.
@@
Several studies have considered the energy aspect of V2X communication, recognising its importance in sustaining large-scale, real-time vehicle connectivity [35, 26, 8, 19, 4]. These works typically explore how different communication control factors can be evaluated to reduce RF energy consumption without compromising reliability. The energy-efficient resource allocation problem in urban V2X scenarios is considered in [35, 26]. It indicates that the appropriate control of ${P}_{\mathrm{t}}$ and interference management are important for energy-efficient V2X. In addition, $\mathrm{SCS}$ and $\mathrm{MCS}$ can also come at the cost of increased power consumption. The research conducted in [8] shows that higher $\mathrm{SCS}$, such as $30$ $\mathrm{kHz}$, can lead to higher energy requirements for accurate transmission and reception due to increased signal bandwidth and processing demands, especially under different traffic conditions. In contrast, it was shown that such low-priority and delay-tolerant applications as traffic monitoring or non-time-sensitive updates, benefit from lower $\mathrm{SCS}$ such as $15$ $\mathrm{kHz}$ [19]. Moreover, lower $\mathrm{MCS}$ levels can conserve energy and minimise resource usage. As demonstrated in [4], selecting a lower $\mathrm{MCS}$ in challenging environments leads to improved transmission efficiency and reduced RF energy consumption without significantly compromising throughput. However, these studies [35, 26, 8, 19, 4] often overlook rural environments, where vehicles are typically more spaced out and signals travel over longer distances with fewer obstructions. Such conditions demand a joint performance evaluation of all three communication control factors to ensure reliable communication while maintaining energy efficiency.
@@
Motivated by these research gaps, this paper investigates the impact of three communication control factors ${P}_{\mathrm{t}}$, $\mathrm{SCS}$, and $\mathrm{MCS}$ on RF energy consumption and communication performance in different rural V2X scenarios. Performance is evaluated using $\mathrm{PRR}$ and the $D_{\mathrm{comm}}$, reflecting the ability to deliver safety messages within safety-critical distances. By analysing $\mathrm{PRR}$ and RF energy consumption under different traffic densities and vehicle speeds, the study aims to provide configuration guidance that ensure reliable, low-latency communication while minimising energy expenditure.
@@
I-B Main Contributions
The main contributions of this paper now can be summarised as follows. Firstly, we investigate the impact of three key communication control factors including $\mathrm{SCS}$, ${P}_{\mathrm{t}}$ of road side unit, and $\mathrm{MCS}$ on safety and RF energy consumption in a connected braking scenario. Our rural scenario represents random layouts in light and heavy traffic conditions, where CAVs exchange time-critical braking messages via 5G NR Sidelink V2X communication. Our second contribution is the development of analytical formulations for the $D_{\mathrm{comm}}$, $\mathrm{PRR}$ and RF energy consumption, which together describe the achievable braking reliability over distance. These formulations provide a comprehensive analysis, linking communication control factors to safety-critical performance indicators. Finally, the impact of communication control factors on the safety and RF energy consumption is validated through extensive simulations, offering valuable insights into the trade-offs between safety and RF energy consumption. Using the analytical results, we identify the appropriate configurations of $\mathrm{SCS}$, ${P}_{\mathrm{t}}$, and $\mathrm{MCS}$ that contributes to an improved safety and energy performance. The analysis forms a baseline for adaptive control strategies on them in future rural CAV safety applications.
The remainder of this paper is organised as follows. Section II introduces the system model and the simulation setup. In Section III, we formulate $D_{\mathrm{comm}}$ and RF energy consumption. In Section IV, we illustrate the simulation results and explore the impact on $\mathrm{PRR}$, $D_{\mathrm{comm}}$ from tuning communication control factors. Section V provides quantitative analysis on the results, and impact of communication control factors on safety and RF energy consumption. Finally, Section VI summarises the main findings of this paper.